OpenStudy (anonymous):

53. What is the simplified form

OpenStudy (anonymous):



OpenStudy (anonymous):

A. 121 b.1/11 c. 11 d 2/242

OpenStudy (hba):

121 is a perfect square.

OpenStudy (anonymous):

Sorry I have a lot of questions I need help with some of them not all. :)

OpenStudy (hba):

like 4 is a perfect square of 2 \[\sqrt{4}=\sqrt{2 \times 2}=2\]

OpenStudy (hba):

Therfore \[\sqrt{121}=\sqrt{{11 \times 11}}=\]

OpenStudy (anonymous):

Okay so then its 2/242

OpenStudy (hba):

No.

OpenStudy (anonymous):

But its doubled

OpenStudy (hba):

No it is not doubled.

OpenStudy (hba):

Sqrt means \[\sqrt{\frac{ 1 }{ 121 }}=(\frac{ 1 }{ 121 })^{1/2}\]

OpenStudy (hba):

\[(\frac{ 1 }{ 11 \times 11 })^{1/2}=(\frac{ 1 }{ 11^2 })^{1/2}\]